OpenStudy (anonymous):

Few Questions. How can I begin? The product of (a − b)(a − b) is a perfect square trinomial. Find the product of (x − 5)2.

OpenStudy (anonymous):

\[ (x-5)^2 \]?

OpenStudy (anonymous):

yes @wio :D

OpenStudy (anonymous):

\[ (x-5)^2=(x-5)(x-5)=x(x-5)-5(x-5) \]Does this make sense?

OpenStudy (anonymous):

 @wio